[Ncep.nhc.nco_contacts] HSU Shift Log for Wed Nov 18 2015 14:42:21 GMT-0500 (EST)

HSU Shift Log michael.j.brennan at noaa.gov
Wed Nov 18 19:42:32 UTC 2015


HSU Shift Log for Wed Nov 18 2015 14:42:21 GMT-0500 (EST)

Shift: H

Forecaster: TBK

Technical issues: Cone graphic for advisory #1 on TD-21E malfunctioned as  
showed a very large area with the track in the upper right corner. EMXI  
track not received in the ATCF despite multiple tries to retrieve it.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20151118/ca16601d/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list